UESTION TWO

The time it takes a randomly selected employee to perform a task is mean value of 120 seconds and standard deviation of 20 seconds.

INCOME-y 28

31

34

30 35 158

            Calculate the probability that a randomly selected employee will complete:

2.1.1 The task within 100 to 130 sec (5)

2.1.2 In more than 75sec


1
Expert's answer
2021-05-17T17:04:11-0400

2.1.1 "P(100<X<130)=P(\frac{100-120}{20}<Z<\frac{130-120}{20})]P(-1<Z<0.5)="

"=P(Z<0.5)-P(Z<-1)=0.6915-0.1587=0.5328."


2.1.2 "P(X>75)=P(Z>\frac{75-120}{20})=P(Z>-2.25)=0.9878."